OpenStudy (mtalhahassan2):

Determine whether the following functions are equivalent. Justify your answer.

OpenStudy (mtalhahassan2):

(x)=x^2+4x−32/ x−4 and g(x)=x(2x−5)−2(x^2−3x−4)

OpenStudy (welshfella):

simplify both functions for the first one first factor x^2 + 4x - 32 and for g(x) use distributive law then simplify like terms.